Find the Taylor series centered at 1 for x3 - 2x? + + 7
Answer #1

Sol 3 وب 2 - 1 = (۶)1 * 2 Taylor Series for f(x) about x= ais = n=0 + دارم . (.. (ا) + (ها} = (م) ( 1 = (۶) در کره (۰) * ا3